New York City is the most expensive city in the United States for lodging. The mean hotel room rate is $203 per night (USA Today, April 30, 2012). Assume that room rates are normally distributed with a standard deviation of $56. Use Table 1 in Appendix B.

a. What is the probability that a hotel room costs $224 or more per night (to 4 decimals)?

b. What is the probability that a hotel room costs less than $141 per night (to 4 decimals)?

c. What is the probability that a hotel room costs between $201 and $300 per night (to 4 decimals)?

d. What is the cost of the 20% most expensive hotel rooms in New York City? Round up to the next dollar.
$  or - Select your answer -morelessItem 5
Answer #1

Part a)

X ~ N ( µ = 203 , σ = 56 )
P ( X >= 224 ) = 1 - P ( X < 224 )
Standardizing the value
Z = ( X - µ ) / σ
Z = ( 224 - 203 ) / 56
Z = 0.375
P ( ( X - µ ) / σ ) > ( 224 - 203 ) / 56 )
P ( Z > 0.375 )
P ( X >= 224 ) = 1 - P ( Z < 0.375 )
P ( X >= 224 ) = 1 - 0.6462
P ( X >= 224 ) = 0.3538

Part b)

X ~ N ( µ = 203 , σ = 56 )
P ( X < 141 )
Standardizing the value
Z = ( X - µ ) / σ
Z = ( 141 - 203 ) / 56
Z = -1.1071
P ( ( X - µ ) / σ ) < ( 141 - 203 ) / 56 )
P ( X < 141 ) = P ( Z < -1.1071 )
P ( X < 141 ) = 0.1341

Part c)

X ~ N ( µ = 203 , σ = 56 )
P ( 201 < X < 300 )
Standardizing the value
Z = ( X - µ ) / σ
Z = ( 201 - 203 ) / 56
Z = -0.0357
Z = ( 300 - 203 ) / 56
Z = 1.7321
P ( -0.04 < Z < 1.73 )
P ( 201 < X < 300 ) = P ( Z < 1.73 ) - P ( Z < -0.04 )
P ( 201 < X < 300 ) = 0.9584 - 0.4858
P ( 201 < X < 300 ) = 0.4726

Part d)

X ~ N ( µ = 203 , σ = 56 )
P ( X > x ) = 1 - P ( X < x ) = 1 - 0.2 = 0.8
To find the value of x
Looking for the probability 0.8 in standard normal table to calculate critical value Z = 0.8416
Z = ( X - µ ) / σ
0.8416 = ( X - 203 ) / 56
X = 250.1296 ≈ $250
P ( X > 250 ) = 0.2